The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 360 mm3, what is the volume of the stack of pennies on the right in cubic millimeters? Use 3.14 for pi. (Hint: only enter numerals in the answer )


1
Expert's answer
2020-07-02T18:11:56-0400

The volume of the stack of pennies on the right is also 360 cubic millimeters.

Proof

Given:- Volume of the stack of pennies on the left = 360 mm^3

& pi = 3.14

now we know that volume = 4/3 pi (r^3)

so

4/3 pi (r^3) = 360 mm^3

r^3 = (360*3)/4 * pi

r^3 = 85.98726115

r = 4.413787008 (radius of a penny)

now for right stack volume :-

volume = 4/3 pi (r^3)

where r = 4.413787008 & pi = 3.14

Volume of the stack of pennies on the right = 4/3 (3.14) ((4.413787008)^3)

= 4/3 (3.14) (85.98726115)

= 360 mm^3

(prooved)